Find the length of the shorter leg of a right triangle if the longer leg is 12 feet more than the shorter leg and the hypotenuse is 12 feet less than twice the shorter leg.

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Let represent the measure of the short leg. Then is the measure of the long leg, and is the measure of the hypotenuse. Use Pythagoras:



Solve for the non-zero root.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Find the length of the shorter leg of a right triangle if the longer leg is 12 feet more than the shorter leg and the hypotenuse is 12 feet less than twice the shorter leg.
----
Let the shorter leg be "x".
Other leg: "x+12"
Hypotenuse: "2x-12"
-------
Equation:
(2x-12)^2 = x^2 + (x+12)^2
-------
4x^2 - 48x + 144 = 2x^2 + 24x + 144
------
2x^2 - 72x = 0
2x(x-36) = 0
Positive solution:
x = 36 (shorter leg)
x+12 = 48 (longer leg)
2x-12 = 60 (hypotenuse)
----------------------
Cheers,
Stan H.